OpenStudy (anonymous):

If you add the digits in a two-digit number and multiply the sum by 7, you get the original number. If you reverse the digits in the two-digit number, the new number is 18 more than the sum of its two digits. What is the original number?

OpenStudy (kitten_is_back):

the digit is x and y If you add the digits in a two-digit number and multiply the sum by 7, you get the original number 7(x+y) = 10x+y 7x+7y = 10x+y 6y = 3x x = 2y If you reverse the digits in the two-digit number, the new number is 18 more than the sum of its two digits 10y+x = x+y+18 9y = 18 y = 2 x = 2(2) = 4 so the original number is 42